# Lecture Notes on High Dimensional Linear Regression

### v1

@@
Alberto Quaini111 Department of Econometrics, Erasmus University of Rotterdam, P.O. Box 1738, 3000 DR Rotterdam, The Netherlands. 222 Please, let me know if you find typos or mistakes at quaini@ese.eur.nl.
(December 20, 2024 (first version: October 17, 2024))
@@ Introduction
At present, the notes focus on linear regression, with an in-depth exploration of the existence, uniqueness, relations, computation, and non-asymptotic properties of the most prominent estimators in this setting: least squares, ridgeless, ridge, and lasso.
@@
It is assumed that readers have a solid background in calculus, linear algebra, convex analysis, and probability theory. Some definitions and results from these fields, relevant for the course, are provided in the Appendix for reference.
@@
  • For a random variable (vector) [matrix] $x$ ($\bm{x}$) [$\bm{X}$], the notation $x\in\mathbb{R}$ ($\bm{x}\in\mathbb{R}^{n}$) [$\bm{X}\in\mathbb{R}^{n\times p}$] means that $x$ ($\bm{x}$) [$\bm{X}$] takes values in $\mathbb{R}$ ($\mathbb{R}^{n}$) [$\mathbb{R}^{n\times p}$].
@@
of explanatory variables $\bm{x}=[x_{1},\ldots,x_{p}]^{\prime}$, where $\bm{\theta}\in\mathbb{R}^{p}$ and $p\in\mathbb{N}$111 An intercept in $f(\bm{x};\bm{\theta})$ can be introduced by adding a constant term to the predictors. The target variable $y$ is also referred to as dependent or output variable, while the explanatory variables $\bm{x}$ are also known as independent variables, predictors or input variables.
@@
$\bm{X}:=\begin{bmatrix}x_{11}&\ldots&x_{1p}\\ \vdots&\ddots&\vdots\\ x_{n1}&\ldots&x_{np}\end{bmatrix}\mathbb{R}^{n\times p}.$
@@ 1.0.1 Estimators and their properties
Let $\bm{M}^{+}$ denote the Moore-Penrose inverse of a generic real-valued matrix $\bm{M}$. We make extensive use of the following useful projections.222 We use the notation $\operatorname{Range}$ ($\operatorname{Ker}$) for the range (kernel) of a matrix. Details on these sets, the Moore-Penrose inverse and orthogonal projections are given in Appendix Section LABEL:app:sec:linear_algebra.
@@
The next proposition demonstrates that, if we fix the design matrix $\bm{X}$, we can focus on regression coefficients in $\operatorname{Range}(\bm{X}^{\prime})$. Indeed, coefficients in this set span all possible linear predictions that can be achieved through $\bm{X}$.
@@
Since an estimator is derived from data, it is a random variable. Intuitively, when comparing two estimators of the same estimand, we prefer the one whose probability distribution is ”more concentrated” around the true value of the estimand. Formally, estimators are compared using several key properties.333 Some of these properties are defined by means of the $l_{2}-$norm. Note that this choice is typical, but arbitrary.
@@
The estimation risk of an estimator $\hat{\bm{\theta}}_{n}$ of $\bm{\theta}_{0}$ measures the difference between the estimate and the estimand as:
@@
The result follows from

### v2

@@
Alberto Quaini [^1] Department of Econometrics, Erasmus University of Rotterdam, P.O. Box 1738, 3000 DR Rotterdam, The Netherlands.  [^2] Please, let me know if you find typos or mistakes at quaini@ese.eur.nl.
(December 3, 2025 (first version: October 17, 2024))
@@ Introduction
At present, the notes focus on linear regression, with an in-depth exploration of the existence, uniqueness, relations, computation, and non-asymptotic properties of prominent estimators in this setting: least squares, ridgeless, ridge, and lasso.
@@
It is assumed that readers have a solid background in mathematical analysis, linear algebra, convex analysis, and probability theory. Some definitions and results from these fields, relevant for the course, are provided in the Appendix for reference.
@@
  • For a random variable (vector) [matrix] $x$ ($\bm{x}$) $\bm{X}$], the notation $x\in\mathbb{R}$ ($\bm{x}\in\mathbb{R}^{n}$) $\bm{X}\in\mathbb{R}^{n\times p}$] means that $x$ ($\bm{x}$) $\bm{X}$] takes values in $\mathbb{R}$ ($\mathbb{R}^{n}$) $\mathbb{R}^{n\times p}$].
@@
of explanatory variables $\bm{x}=[x_{1},\ldots,x_{p}]^{\prime}$, where $\bm{\theta}\in\mathbb{R}^{p}$ and $p\in\mathbb{N}$. [^1] An intercept in $f(\bm{x};\bm{\theta})$ can be introduced by adding a constant term to the predictors. The target variable $y$ is also referred to as dependent or output variable, while the explanatory variables $\bm{x}$ are also known as independent variables, predictors or input variables.
@@
$\bm{X}:=\begin{bmatrix}x_{11}&\ldots&x_{1p}\\ \vdots&\ddots&\vdots\\ x_{n1}&\ldots&x_{np}\end{bmatrix}\in\mathbb{R}^{n\times p}.$
@@ 1.0.1 Estimators and their properties
Let $\bm{M}^{+}$ denote the Moore-Penrose inverse of a generic real-valued matrix $\bm{M}$. We make extensive use of the following useful projections. [^2] We use the notation $\operatorname{Range}$ ($\operatorname{Ker}$) for the range (kernel) of a matrix. Details on these sets, the Moore-Penrose inverse and orthogonal projections are given in Appendix Section LABEL:app:sec:linear_algebra.
@@
The next proposition demonstrates that, if we fix the design matrix $\bm{X}$, we can focus on regression coefficients in $\operatorname{Range}(\bm{X}^{\prime})$. Indeed,
$\mathbb{R}^{p}=\operatorname{Ker}(\bm{X})+\operatorname{Range}(\bm{X}^{\prime})$
and coefficients in $\operatorname{Ker}(\bm{X})$ are immaterial for linear prediction, while coefficients in $\operatorname{Range}(\bm{X}^{\prime})$ span all possible linear predictions that can be achieved through $\bm{X}$.
@@
Since an estimator is derived from data, it is a random variable. Intuitively, when comparing two estimators of the same estimand, we prefer the one whose probability distribution is ”more concentrated” around the true value of the estimand. Formally, estimators are compared using several key properties. [^3] Some of these properties are defined by means of the $l_{2}$-norm. Note that this choice is typical, but arbitrary.
@@
The estimation risk of an estimator $\hat{\bm{\theta}}_{n}$ of $\bm{\theta}_{0}$ measures the difference between the estimator and the estimand as:
@@
While the MSE and and the MPR are deterministic quantities, the estimation risk and the predictive risk are stochastic, unless we treat the design matrix as fixed.
@@
The result follows from $\mathbb{E}[\hat{\bm{\theta}}_{n}-\mathbb{E}[\hat{\bm{\theta}}_{n}]]=0$ and